# BOP Challenge 2024 on Model-Based and Model-Free 6D Object Pose Estimation

### v1

@@ 3.3 Pre-training dataset
As in 2023, we provided over 2M training images showing 50K+ diverse objects, which are not included in BOP-Classic nor BOP-H3 and can be therefore used for pre-training methods for tasks on unseen objects. The images were originally synthesized for MegaPose [23] using BlenderProc [7, 6, 8]. The objects are from the Google Scanned Objects [10] and ShapeNetCore [4] datasets.
@@ 4 Results and discussion
4.1 Experimental setup
Participants were submitting results to the online evaluation system at bop.felk.cvut.cz from May 29th, 2024 until the deadline on Nov 29th, 2024.333Evaluation scripts are in: github.com/thodan/bop_toolkit A method had to use a fixed set of hyper-parameters across all objects and datasets. In the model-based tasks, a method could render images of the 3D models or use a subset of the BlenderProc images originally provided for BOP 2020 [19] – the method could use as many images from this set as could be rendered within the limited onboarding time (rendering and any additional processing had to fit within 5 minutes, considering that rendering of one BlenderProc image takes 2 seconds). In the model-free tasks, a method could apply any reconstruction method to the onboarding videos to obtain 3D models of test objects, but the whole onboarding of an object had to fit within 5 minutes.
Not a single pixel of test images may have been used for training and onboarding, nor the individual GT that are publicly available for test images of some datasets. Ranges of the azimuth and elevation camera angles, and a range of the camera-object distances determined by the GT poses from test images are the only information about the test set that may have been used during training and onboarding. Only subsets of test images were used (see Tab. 1) to remove redundancies and speed up the evaluation, and only object instances for which at least of the projected surface area is visible were considered in the evaluation.
@@ 4.1 Experimental setup
| 21 | GenFlow-MultiHypo16 [29] | 2023 | CNOS-FastSAM | GenFlow | RGB-D | PBR | RGB-D | 63.5 | 52.1 | 86.2 | 53.4 | 55.4 | 77.9 | 83.3 | 67.4 | 34.58 | |
| 22 | GenFlow-MultiHypo [29] | 2023 | CNOS-FastSAM | GenFlow | RGB | PBR | RGB-D | 62.2 | 50.9 | 84.9 | 52.4 | 54.4 | 77.0 | 81.8 | 66.2 | 21.46 | |
@@
4.2 Challenge tracks
Participants competed on seven challenge tracks, each defined by object onboarding setup (model-based, model-free), a task (6D localization, 6D detection, 2D detection), and dataset group (BOP-Classic-Core, BOP-H3):
Track 1: Model-based 6D localization on BOP-Classic-Core
Track 2: Model-based 6D detection on BOP-Classic-Core
Track 3: Model-based 2D detection on BOP-Classic-Core
Track 4: Model-based 6D detection on BOP-H3
Track 5: Model-based 2D detection on BOP-H3
Track 6: Model-free 6D detection on BOP-H3
Track 7: Model-free 2D detection on BOP-H3
4.3 Model-based 6D loc. on BOP-Classic-Core
Results on model-based 6D object localization of unseen objects in BOP-Classic-Core are shown in Table 2. Among the 44 new methods submitted in 2024, twenty outperform GenFlow [29] (row #21), the best from the 2023 challenge. The top-performing method, FreeZeV2.1 [3], achieves 82.1 AR—22% higher than GenFlow, with 28% faster runtime. It is also only 4% ARC behind GPose2023, the best method for seen objects. FreeZeV2.1 combines 2D features from frozen DINOv2 [34] and 3D features from GeDi [36], and uses detections from SAM6D [24], NIDS [27], and CNOS [32]. It renders 162 templates per object, extracts 2D visual features, backprojects them to 3D, aggregates, and performs 3D-3D matching, followed by ICP and symmetry-aware refinement.
Another notable method is FRTPose.v1, ranked second with 81.8 AR. During on-boarding, a ResNet34 is trained to predict 3D surface coordinates from 2D inputs and object masks, taking about 4.5 minutes per object on an H100 GPU. It then uses PnP for initial pose estimation and refines it using FoundationPose [44]. As all methods use a two-stage approach—first detecting objects in 2D, then estimating the 6D pose per detection—we analyzed the qualitative results and found that the most common cause for missing or inaccurate 6D pose estimates are erroneous 2D detections.
@@ 4.2 Challenge tracks
4.4 Model-based 6D det. on BOP-Classic-Core
Table 3 presents the results of the 20 entries on model-based 6D detection of unseen objects. While the ranking among different methods is consistent between Track 1 and Track 2: FreeZeV2.1 is the best method, followed by FRTPose.v2 and Co-op [30]; the relative improvement of FreeZeV2.1 over Co-op is more significant in 6D detection with 18.8% (#1 and #7 of Tab. 3) than in 6D localization with only 6.5% (#1 and #8 of Tab. 2).
Interestingly, Co-op (row #7 in Table 3) shows a different trend from Track 1, with coarse results outperforming refined ones (e.g., row #7 over #8, and #9 over #10 and #11). Co-op uses a three-stage pipeline: coarse estimation with pretrained CrocoV2 [42], refinement via optical flow (similar to GenFlow [29] but without RAFT [39]), and scoring with a MegaPose-style [23] network. The coarse results outperforming refined ones can be attributed to the scoring network being trained only on template-crop input pairs of the same object, making it unreliable when the template and crop candidates often display different objects.
4.5 Model-based 2D det. on BOP-Classic-Core
Tab. 4 shows the results of model-based 2D detection for unseen objects on BOP-Classic-Core datasets (Track 3). The best method, MUSE, uses a novel similarity metric based on both class and patch embeddings and weights the contributions of 3D template views by a von Mises Fisher distribution. MUSE achieves 52.0 AP, + 21% more accurate compared to the best method from the 2023 challenge, CNOS [32] (#10 in Table 4). However, MUSE is still -53% behind the best method for seen objects (GDet2023).
4.6 Model-based 6D detection on BOP-H3
Table 5 presents results for model-based 6D detection of unseen objects on BOP-H3 (Track 4), with 4 entries. All methods used RGB/monochrome images, as depth is unavailable for HOT3D and HANDAL. GigaPose and GigaPose+GenFlow achieve 9.4 and 31.2 AP on BOP-H3, compared to 12.3 and 50.4 AP on BOP-Classic-Core (Track 2, rows #20 and #18 in Table 3). OPFormer, a new method, aggregates DINOv2 [34] template descriptors using a transformer with 3D positional embeddings and predicts poses via RANSAC-PnP. On HOPEv2, OPFormer outperforms GigaPose by 110% but is 5% less accurate than GigaPose+GenFlow after refinement.
4.7 Model-based 2D detection on BOP-H3
Tab. 6 shows the results of model-based 2D detection for unseen objects, on BOP-H3 datasets. Similar to the results on BOP-Classic-Core datasets, MUSE outperforms CNOS [32] while the relative improvement is not the same: MUSE is 21.5% more accurate than CNOS in BOP-Classic-Core (#1, #10 in Table 4) while the improvement in BOP-H3 is 33.1% (#1, #2 in Table 6).
4.8 Model-free 2D detection on BOP-H3
The new task of model-free 2D detection of unseen objects on the BOP-H3 datasets received 3 entries, as presented in Tab. 7. GFreeDet [26] introduces a model-free 2D detection method that leverages Gaussian splatting [22] to reconstruct objects, then use DINOv2 [34] for template matching. Additionally, on HOPEv2 dataset, GFreeDet is 5% more accurate than CNOS [32] (36.4 vs 34.5 AP) as shown in the leaderboard for this track444 bop.felk.cvut.cz/leaderboards.
@@ 5 Awards
FreeZeV2.1 [3] by Andrea Caraffa, Davide Boscaini, Amir Hamza, and Fabio Poiesi
Co-op [30] by Sungphill Moon and Hyeontae Son
FoundationPose [44] by Bowen Wen, Wei Yang, Jan Kautz, and Stan Birchfield
FRTPose.v1 (anonymous for now)
GigaPose [33] by Van Nguyen Nguyen, Thibault Groueix, Mathieu Salzmann, and Vincent Lepetit
MUSE (anonymous for now)
SAM6D [24] by Jiehong Lin, Lihua Liu, Dekun Lu, and Kui Jia
GenFlow [29] by Sungphill Moon, Hyeontae Son, Dongcheol Hur, and Sangwook Kim
CNOS [32] by Van Nguyen Nguyen, Thibault Groueix, Georgy Ponimatkin, Vincent Lepetit, and Tomas Hodan
GFreeDet [26] by Xingyu Liu, Yingyue Li, Chengxi Li, Gu Wang, Chenyangguang Zhang, Ziqin Huang, and Xiangyang Ji
@@ 6 Conclusions
In BOP 2024, methods for 6D localization of unseen objects have almost reached the accuracy of their counterparts for seen objects. While some of the methods for unseen objects now take less than 1 second, further improvements are required to support real-time applications. We hope that the model-free tasks introduced in the 2024 challenge will unlock new types of applications, making object pose estimation significantly more practical. In BOP’25, we are introducing BOP-Industrial datasets and the multi-view problem setup to represent real-world applications in industrial robotics. The evaluation system at bop.felk.cvut.cz stays open and raw results of all methods are publicly available.

### v4

@@ 2.4 Synthetic training dataset
As in 2023, we provided over 2M synthetic training images showing 50K+ diverse objects from the GSO [10] and ShapeNetCore [4] datasets. These objects are not included in BOP-Classic nor BOP-H3 and the dataset can be therefore used for training methods for tasks on unseen objects. The images were originally synthesized for MegaPose [25] using BlenderProc [7, 6, 8].
3 Challenge tracks
In 2024, participants competed on seven challenge tracks on unseen objects (target objects are not seen during training), with each track defined by a task (6D localization, 6D detection, 2D detection), type of object onboarding (model-based, model-free), and dataset group (BOP-Classic-Core, BOP-H3):
Track 1: Model-based 6D localization on BOP-Classic-Core
Track 2: Model-based 6D detection on BOP-Classic-Core
Track 3: Model-based 2D detection on BOP-Classic-Core
Track 4: Model-based 6D detection on BOP-H3
Track 5: Model-based 2D detection on BOP-H3
Track 6: Model-free 6D detection on BOP-H3
Track 7: Model-free 2D detection on BOP-H3
We primarily focused on the more practical detection tasks (Track 2-7) while keeping the classical model-based 6D localization on BOP-Classic-Core to enable direct comparison with previous years (Track 1).
On all tracks, methods go through three phases: (1) training, where methods can be trained on non-target objects, (2) object onboarding, where methods have to quickly onboard target objects, and (3) inference, where methods are expected to produce predictions about target objects seen in test images. The following sections describe these phases in more detail.
3.1 Training phase
At training time, methods can be trained on the provided dataset of 2M synthetic images, originally synthesized using BlenderProc [7, 6] for MegaPose [25]. The images show training (non-target) objects annotated with ground-truth 6D poses, modal 2D segmentation masks (covering the visible object parts), and amodal 2D bounding boxes (covering the whole object silhouette, including the occluded parts). The dataset also includes color 3D mesh models of the objects. Methods are also free to use other datasets as long as the images did not show target objects.
3.2 Object onboarding phase
After training, methods can spend up to 5 minutes of the wall-clock time on a single computer with one GPU to onboard a target object.111The time is measured from the point right after the raw data (a 3D mesh model in the model-based setup and reference video(s) in the model-free setup) is loaded to the point when the object is onboarded. The generated representation of onboarded objects (object templates, neural radiance fields, etc.) need to be fixed after onboarding (cannot be updated based on test images).
Model-based onboarding: Methods are provided 3D models of target objects that are not seen at training. To onboard an object, methods can, e.g., fine-tune a neural network or render images of the 3D models, but cannot use any real images of the objects.
Model-free onboarding: Methods are provided video(s) of target objects that are not seen at training. 3D models of the objects are not available. Methods can use either static or dynamic onboarding videos (Fig. 3, Sec. 2.2), e.g., to reconstruct a 3D model, render novel views, or fine-tune a neural network.
3.3 Inference phase
Input: Methods are given a real-world test image showing an arbitrary number of instances of an arbitrary number of target objects, with all objects being from one specified dataset (e.g., YCB-V). Depending on the dataset, the test image may be RGB-D, RGB, or monochrome. In case of 6D localization, methods are additionally provided identities of objects visible in the test image in the form of a list , where is the number of instances of object for which at least of the projected surface area is visible in the test image. No prior information about the visible object instances is provided in the detection tasks.
As most recent methods for 6D object localization and detection start with a 2D detection stage, participants in the 6D localization and detection tasks are also encouraged to evaluate their methods using default 2D detections/segmentations produced by CNOS [35]. Starting from the same 2D detections enables direct comparison of the object pose estimation stages.
Output: For each test image, methods produce predictions with confidence scores and run time.222Although the run time is measured on user machines, which may be different for each method, it still provides a general sense of how fast each method is. The time is measured from the point right after the raw data (the image, 3D models, etc.) is loaded to the point when predictions for all objects in the image are available (including the time needed to generate default 2D detections if used). Predictions are in the form of amodal 2D bounding boxes for 2D detection and in the form of 6D object poses for 6D localization and detection. A 6D object pose is defined by a matrix , where is a 3D rotation matrix, and is a 3D translation vector from the 3D object model space to the 3D camera space. In all tasks, methods need to produce predictions only for object instances for which at least of the projected surface area is visible in the given test image.
3.4 Evaluation methodology
Measuring error of 6D object poses. The error of an estimated pose w.r.t. the ground-truth pose can be calculated by the following pose-error functions (see Sec. 2.2 of [20] for details): (1) MSSD (Maximum Symmetry-Aware Surface Distance) which considers a set of pre-identified global object symmetries and measures the surface deviation in 3D, (2) MSPD (Maximum Symmetry-Aware Projection Distance) which considers the object symmetries and measures the perceivable deviation, (3) VSD (Visible Surface Discrepancy) which treats indistinguishable poses as equivalent by considering only the visible object part. An estimated pose is considered correct w.r.t. a pose-error function , if , where and is a threshold of correctness.
Evaluating 6D object localization. The fraction of object instances for which a correct pose is estimated is called Recall. The Average Recall w.r.t. a pose-error function on a dataset , denoted as , is defined as the average of the Recall rates calculated for multiple settings of and, in the case of VSD, also for multiple settings of the misalignment tolerance . The accuracy on a dataset is measured by: , and the overall accuracy, denoted as AR, is defined as the average of scores over selected datasets.333When calculating AR, scores are not averaged over objects before averaging over datasets, which is done when calculating AP (in 2D/6D detection) to comply with the original COCO evaluation methodology [27].
Evaluating 2D object detection. The detection accuracy is measured by the Average Precision (AP, also known as mAP), following the evaluation methodology from the COCO 2020 Object Detection Challenge [27]. Specifically, a per-object score is calculated by averaging the precision values at multiple thresholds on the Intersection over Union (IoU) of 2D bounding boxes: . The accuracy of a method on a dataset is measured by calculated by averaging per-object scores over objects from the dataset. The overall accuracy, denoted as AP, is defined as the average of the per-dataset scores.444Up to most confident detections per image are considered. Correct detections for instances visible from less than are not counted as false positives.
Evaluating 6D object detection. The detection accuracy is measured by the Average Precision (AP), similarly as for 2D object detection but using pose-error functions instead of the IoU of 2D bounding boxes. Specifically, for each pose-error function 555The VSD pose-error function is not considered as it is more expensive to calculate (6D detection requires evaluating more pose estimates than 6D localization) and needs depth images which are not available in HOT3D and HANDAL., the per-object accuracy is defined by averaging the precision values calculated at multiple thresholds of correctness (defined in Sec. 2.4 of [20]). The accuracy on a dataset w.r.t. a pose-error function is calculated by averaging per-object scores over objects from the dataset. The accuracy on the dataset is then defined as , and the overall accuracy, denoted as AP, is defined as the average of scores over selected datasets.4
@@ 4 Results and discussion
| 21 | GenFlow-MultiHypo16 [32] | 2023 | CNOS-FastSAM | GenFlow | RGB-D | PBR | RGB-D | 63.5 | 52.1 | 86.2 | 53.4 | 55.4 | 77.9 | 83.3 | 67.4 | 34.6 | |
| 22 | GenFlow-MultiHypo [32] | 2023 | CNOS-FastSAM | GenFlow | RGB | PBR | RGB-D | 62.2 | 50.9 | 84.9 | 52.4 | 54.4 | 77.0 | 81.8 | 66.2 | 21.5 | |
@@
Participants were submitting results to the online evaluation system from May 29th until November 29th, 2024. We received submissions for all but Track 6, presumably due to the limited time (BOP-H3 datasets were released later) and the extra effort required to prepare model-free methods. In this section we aim to provide a high-level summary of the results and refer the reader to the online system for more per-submission details. Evaluation scripts used by the system are publicly available in the BOP toolkit.
@@
4.1 Experimental setup
A method had to use a fixed set of hyperparameters across all objects and datasets. In the model-based onboarding, a method could render images of the 3D models or use a subset of the BlenderProc images (originally provided for BOP 2020 [20]), assuming that rendering a single BlenderProc image takes 2 seconds. The onboarding phase had to be done within 5 minutes on a single GPU. Note that methods marked as RGB-only could only use RGB/monochrome channels for onboarding and inference. Not a single pixel of test images may have been used for training and onboarding, nor ground-truth annotations that are available for test images of some datasets. Ranges of the azimuth and elevation camera angles, and the range of camera-object distances determined by the ground-truth poses from test images are the only information about the test set that may have been used during training and onboarding. Only subsets of test images were used (see Tab. 1) to remove redundancies and speed up the evaluation.
4.2 Model-based 6D localization of unseen objects
Among the 44 new entries from 2024 for model-based 6D object localization on BOP-Classic-Core (Tab. 2), 20 entries (19 RGB-D, 1 RGB) outperform the best method from 2023, an RGB-D variant of GenFlow [32] (#21 in Tab. 2). The best method from 2024, FreeZeV2.1 [3], is 22% more accurate (82.1 vs. 67.4 AR) and 39% faster than GenFlow. Notably, FreeZeV2.1 is only 3.5 AR behind GPose2023, currently the best method for seen objects.
FreeZeV2.1 is an RGB-D method which, in the winning variant, relies on 2D segmentation masks from three methods: SAM6D [26], NIDS [30], and CNOS [35]. FreeZeV2.1 estimates a 6D pose with a confidence score for each mask and keeps only the best estimates by applying non-maximum suppression. Given a segmentation mask, the pose is estimated in a RANSAC-fitting scheme from 3D-3D correspondences established by matching features (2D visual features from DINOv2 [37] concatenated with 3D geometry features from GeDi [39]) between the query image and the 3D model (the features are extracted from renderings of the model and aggregated on its surface). Pose estimates are further refined by depth-based ICP and symmetry-aware refinement.
Another notable method is FRTPose.v1, ranked 2nd, which quickly trains a ResNet34 network (4.3 min per object on RTX 4090) at onboarding. Poses are estimated from predicted dense 2D-3D correspondences and refined by FoundationPose [47].
A more practical method is Co-op [33] (#9 in Tab. 2), which is 25 times faster (0.8 s per image) and 13% more accurate than GenFlow [32]. Co-op estimates poses with a template-based approach, followed by a flow-based refinement similar to GenFlow.
@@ 4.2 Model-based 6D localization of unseen objects
4.3 Model-based 6D detection of unseen objects
BOP-Classic-Core (Tab. 3). Ranking of methods on Track 2 is consistent with Track 1, suggesting that the 6D detection and 6D localization tasks present similar difficulties and that learnings from previous BOP challenges, focused on the 6D localization task, are also relevant for the newly introduced and more practical 6D detection task. On both tracks, FreeZeV2.1 is the best method, although the improvement of FreeZeV2.1 over Co-op is more significant on Track 2 than on Track 1: 18.8% (#1 and #4 in Tab. 3) vs. only 6.5% (#1 and #8 in Tab. 2). Adding refinement to Co-op tends to decrease the 6D detection score (see coarse poses on #6 in Tab. 3 vs. #7 with single- and #8 with multi-hypotheses refinement). This is a surprising result that is worth investigating.
BOP-H3 (Tab. 5). The best entry is GigaPose [36], a coarse pose estimation method using 2D-2D correspondences established between the input image and pre-rendered templates, followed by the GenFlow [32] refinement. GigaPose and GigaPose+GenFlow achieve 12.3 and 50.4 AP on BOP-Classic-Core (#17 and #15 in Tab. 3), while only 9.4 and 31.2 AP on BOP-H3, suggesting that BOP-H3 is more difficult. Another method on this track is OPFormer, which follows a similar pipeline as FoundPose [38], but predicts correspondences using a trained transformer head instead of relying on a simple kNN matching of DINOv2 features [37].
4.4 Model-based 2D detection of unseen objects
BOP-Classic-Core (Tab. 4). The best 2024 method, MUSE, follows a similar pipeline as the best 2023 method, CNOS [35]: detection proposals are generated using SAM2 [40] and Grounding DINO [28] and then matched to templates using DINOv2 [37]. The key difference is in the matching stage, where MUSE introduces a novel similarity metric that leverages both class and patch tokens, whereas CNOS only uses class tokens. MUSE is 21% more accurate than CNOS (#1 vs. #10 in Tab. 4), however, MUSE is still -35% behind the best method for seen objects (GDet2023).
4.5 Model-free 2D detection of unseen objects
The only entry, GFreeDet [29], reconstructs a 3DGS model [23] from onboarding videos and renders templates from equidistantly sampled viewpoints. At inference, GFreeDet uses SAM [24] to generate proposals and DINOv2 [37] for proposal-template matching. Notably, GFreeDet outperforms the model-based CNOS on HOPEv2 and HANDAL (#1 in Tab. 7 vs. #2 Tab. 6).
@@ 5 Awards
Authors of awarded entries: FreeZeV2.1 [3] by Andrea Caraffa, Davide Boscaini, Amir Hamza, Fabio Poiesi; Co-op [33] and GenFlow [32] by Sungphill Moon, Hyeontae Son, Dongcheol Hur, Sangwook Kim; FoundationPose [47] by Bowen Wen, Wei Yang, Jan Kautz, Stan Birchfield; GigaPose [36] by Van Nguyen Nguyen, Thibault Groueix, Mathieu Salzmann, Vincent Lepetit; SAM6D [26] by Jiehong Lin, Lihua Liu, Dekun Lu, Kui Jia; CNOS [35] by Van Nguyen Nguyen, Thibault Groueix, Georgy Ponimatkin, Vincent Lepetit, Tomas Hodan; GFreeDet [29] by Xingyu Liu, Yingyue Li, Chengxi Li, Gu Wang, Chenyangguang Zhang, Ziqin Huang, Xiangyang Ji; FRTPose.v1 and MUSE are anonymous for now.
@@ 6 Conclusions
In 2024, methods for 6D localization of unseen objects have almost reached the accuracy of methods for seen objects. While some of these methods now take less than 1 second per image, further speed up is needed for real-time applications. Rankings on 6D localization and 6D detection are consistent, suggesting that learnings from past BOP challenges are relevant also for the newly introduced and more practical 6D detection task. 2D detection of unseen objects has improved significantly, but is still noticeably behind 2D detection of seen objects. The evaluation system stays open and awaits new submissions on the model-free tasks.